OpenStudy (anonymous):

Show work What NUMBER is 42% of 150 ?

OpenStudy (lilg132):

remember to click good answer :D

OpenStudy (farmdawgnation):

Ehh, lilg, I may just be having a brain fail - but why would this not be 0.42(150)?

OpenStudy (lilg132):

that will give you 45% of 150

OpenStudy (anonymous):

150 - 100% x - 42% x=42*150/100

OpenStudy (lilg132):

oops

OpenStudy (lilg132):

i misread the question

OpenStudy (anonymous):

so ask farmdawgnation to remove 10medals from you

OpenStudy (lilg132):

please do

OpenStudy (anonymous):

Do i even have any medals

OpenStudy (lilg132):

u do now

OpenStudy (anonymous):

Thx :)

OpenStudy (anonymous):

How do i solve this problem though... What Number is 42% of 150?

OpenStudy (lilg132):

the way i do it is 150 / 100 = 1.5( this is 1% of 150) then multiply that number by the percentage you want to find which is 42 1.5 * 42 = ??

OpenStudy (s):

For this type of problems think about it this way. 100% is the whole thing. 50% is a half. So 42% would be something under a half of your number. Your number is 150. 42% is the same as 0.42 (or 42%/100=0.42). So you just multiply 150*0.42=63.

OpenStudy (lilg132):

i thought i read what number is it if 45% is 150 :p

OpenStudy (lilg132):

*42%

OpenStudy (anonymous):

or use proportion as I did

OpenStudy (anonymous):

so the anwser is...................................................

OpenStudy (farmdawgnation):

See S's answer.